OpenStudy (anonymous):

solve for x: log x + log (x-3)=1 3√x-2 -8=8 12x^2=3x 4e^2x=12

OpenStudy (anonymous):

log(x/(x-3))=log 10 x=10x-30 9x=30 x=10/3

OpenStudy (anonymous):

12x^2=3x 3x(4x-1)=0 x=0,x=1/4